# Bifurcations and canards in the FitzHugh-Nagumo system: a tutorial in fast-slow dynamics

### v1

@@ 5. Bifurcations
Figure 7. Variation of the real and imaginary parts of the eigenvalues of the Jacobian matrix of the system (23), evaluated at the equilibrium point, as function of the parameter $b$ ($\lambda_{\varepsilon}^{+}$ with solid line, $\lambda_{\varepsilon}^{-}$ with dashed line, and Hopf bifurcation point) for different values of $\varepsilon$, $\varepsilon=1$ (green), $\varepsilon=0.5$ (yellow), $\varepsilon=0.1$ (blue) e $\varepsilon=0$ (violet).
Figure 8. (a) - Bifurcation of the equilibria of the system (22) as a function of the parameter $b$, with $\varepsilon=0.5$. Equilibrium point $\mathrm{E}_{0}$ is always unstable (dashed line). Equilibria $\mathrm{E}_{-}$ and $\mathrm{E}_{+}$ are unstable for $1/4<b<b_{0.5}^{H}\approx 0.3666$ and stable $b>b_{0.5}^{H}$ (solid line). Bifurcation points at $b=1/4$ (Pitchfork) and at $b=b_{0.5}^{H}$ (supercritical Hopf). (b) Amplitude of the periodic solution around the point $\mathrm{E}_{+}$ as a function of $b$. Supercritical Hopf bifurcation (black dot) leads to an unstable periodic solution that collides with the homoclinic orbit generated by the saddle point $\mathrm{E}_{0}$. Homoclinic bifurcation at $b_{0.5}^{Hom}\approx 0.36932$ (blue dot). (c) - Example of unstable periodic solutions of the system (23) generated by the Hopf bifurcation, for $b_{0.5}^{H}\leq b\leq b_{0.5}^{Hom}$ and $\varepsilon=0.5$. Homoclinic orbit (black line). $\mathcal{C}_{0}$ and $\mathrm{E}_{0}$ in red, vertical lines $x=x\left(b_{0.5}^{H}\right)$ (yellow) and $x=x\left(b_{0.5}^{Hom}\right)$ (black). Colour scale for the variation of the parameter $b$.
@@ 6. Singularities and Canards
As we have seen in case (i) of Section 5, when the system’s equilibrium is also a fold point, a subcritical Hopf bifurcation occurs, giving rise to a stable periodic solution. In Sections 3 and 4.1, we found another periodic solution which is not related to the Hopf bifurcation. Thus, for different parameters, we have two stable periodic solutions with distinct dynamical properties. In this section, the canard phenomena bridges the local periodic solution from the Hopf bifurcation to the global limit cycle.
As illustrated in Figure 6(c), the existence in (22) of a periodic solution around the equilibrium point implies that the orbit traverses a part of its trajectory close to the repelling region of the critical manifold $\mathcal{C}_{0}$. This behaviour seems to be contradictory since we found a solution that is close to the unstable branch of $\mathcal{C}_{0}$.
@@
Given a singular $(1,1)$–fast–slow system parametrised by $\lambda\in{\mathbb{R}}$. Let $\mathfrak{p}=\left(x,y\right)$ be a fold point. The point $\mathfrak{p}$ is called a singular fold if:
@@
Theorem 4 (Krupa & Szmolyan, 2001a and 2001b [14, 15]. See also Theorems 8.1.3 and 8.2.1 of [16]).
@@
$\lambda_{c}\left(\varepsilon\right)=-\left(B+A\right)\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
@@
$\lambda_{\mathrm{H}}\left(\varepsilon\right)=-B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
@@
An important consideration is the empirical difficulty in finding canards, since the smaller $\varepsilon$ is, the narrower the interval of $\lambda$ values for which canards appear. This small interval separates the $\lambda$ values at which a limit cycle occurs from those where it either undergoes a Hopf bifurcation or does not have any periodic solution. For this reason, the dynamics produced in the system around this small interval is called the canard explosion.
@@
$\bar{\mathrm{c}}_{\mathrm{H}}\left(\varepsilon\right)=0\cdot\,\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right)$
@@
$\bar{\mathrm{c}}_{c}\left(\varepsilon\right)=-\dfrac{3}{8}\,\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).\qquad\diamond$
Figure 10 illustrates the emergence of canards in system (22) with $\varepsilon=0.5$ and $\varepsilon=0.1$. As we may see, the stable periodic solutions emerge from the subcritical Hopf bifurcation for $c=2/\sqrt{3}$ (analogous for $c=-2/\sqrt{3}$). As mentioned above, some of these periodic solutions remain close to the repelling region of the critical manifold, making them canards. In Figures 10(a) and 10(c), two types of orbits can be distinguished: the first type, coloured in yellow, consists of slow trajectories near $\mathcal{C}_{0\text{M}}$ and $\mathcal{C}_{0\text{R}}$ and one fast trajectory; the second type, coloured in green, which surrounds the other fold point, consists of three slow trajectories near $\mathcal{C}_{0\text{L}}$, $\mathcal{C}_{0\text{M}}$ and $\mathcal{C}_{0\text{R}}$ and two fast trajectories. We refer to these two types of canards as headless canard and headed canard, respectively [^2] This terminology is motivated by the literal meaning of the word ”canard” in French (canard = duck). In this system, it is a bit more complicated to understand why, but if the reader tilts their head to the left they will see that the canard orbits resemble the shape of a duck, with or without head..
Note that the amplitude of the canards grows almost instantaneously. As stated in [16], this was expected, but it is still surprising. The nearly vertical lines observed in Figures 10(b) and 10(d) show canards with amplitudes, $\mathcal{A}\left(c\right)\in\left[2,20\right]$ for $c\approx 1.150077$ in the case $\varepsilon=0.5$ and for $c\approx 1.153794$ in the case $\varepsilon=0.1$.
Figure 10. (a) - Flow of the system (22), for $\varepsilon=0.5$. Critical manifold in red, attracting regions (solid line) and repelling region (dashed line), fold points (black), canards without head (yellow), canards with head (green), and limit cycles (blue). (b) - Bifurcation diagram of the system observed in (a). (c) - System (22), with $\varepsilon=0.1$. (d) - Bifurcation diagram of the system observed in (c).
@@ 7. Discussion and Further Work
This article studies the dynamics of a fast-slow system derived from the FitzHugh-Nagumo model according to the geometric singular perturbation theory and the bifurcation theory methods. In particular, we provide an analytic proof that the fast-slow FH-N system presents relaxation oscillation dynamics as well as periodic solutions induced by Hopf bifurcation. We emphasise the emergence of a homoclinic orbit and canards connecting these two distinct dynamics. We illustrate each result with numerical computations.
When coupling two FH-N systems through the slow equations, we have performed numerical simulations that suggest the existence of mixed-mode oscillations where the dynamics is characterised by the existence of orbits that alternate periodically between large and small amplitude oscillations [6]. Recently, Kristiansen & Pedersen [13] showed that the coupling through the fast equations of two distinct FH-N models produces mixed-mode oscillations induced by a cusp singularity present in this system. We believe that a similar result may explain our numerical findings.
@@
$\phantom{\varepsilon\,}\dfrac{\mathrm{d}\theta}{\mathrm{d}t}=\omega$
@@ Acknowledgments
The first and second authors have partially supported by CMUP (UID/MAT/00144/2020), which is funded by FCT with national (MCTES) and European structural funds through the programs FEDER, under the partnership agreement PT2020. The third author has been supported by the Project CEMAPRE/REM – UIDB/05069/2020 financed by FCT/MCTES through national funds.

### v3

@@ 5. Bifurcations
Figure 7. Evolution of the real and imaginary parts of the eigenvalues of the Jacobian matrix of the system (24), evaluated at the equilibrium point, as function of the parameter $b$ ($\lambda_{\varepsilon}^{+}$ with solid line, $\lambda_{\varepsilon}^{-}$ with dashed line, and Hopf bifurcation point) for different values of $\varepsilon$, $\varepsilon=1$ (green), $\varepsilon=0.5$ (yellow), $\varepsilon=0.1$ (blue) e $\varepsilon=0$ (violet).
Figure 8. (a) - Bifurcation of the equilibria of the system (24) as a function of the parameter $b$, with $\varepsilon=0.5$. Equilibrium point $\mathrm{E}_{0}$ is always unstable (dashed line). Equilibria $\mathrm{E}_{-}$ and $\mathrm{E}_{+}$ are unstable for $1/4<b<b_{0.5}^{H}\approx 0.3666$ and stable $b>b_{0.5}^{H}$ (solid line). Bifurcation points at $b=1/4$ (Pitchfork) and at $b=b_{0.5}^{H}$ (supercritical Hopf). (b) Length of the periodic solution around the point $\mathrm{E}_{+}$ as a function of $b$. Supercritical Hopf bifurcation (black dot) leading to an unstable periodic solution that collides with the homoclinic orbit associated to the saddle point $\mathrm{E}_{0}$. Homoclinic bifurcation at $b_{0.5}^{Hom}\approx 0.36932$ (blue dot). (c) - Example of unstable periodic solutions of the system (24) generated by the Hopf bifurcation, for $b_{0.5}^{H}\leq b\leq b_{0.5}^{Hom}$ and $\varepsilon=0.5$. Homoclinic orbit (black line). Critical manifold $\mathcal{C}_{0}$ and equilibrium $\mathrm{E}_{0}$ in red, vertical lines $x=x\left(b_{0.5}^{H}\right)$ (yellow) and $x=x\left(b_{0.5}^{Hom}\right)$ (black). Colour scale for the variation of the parameter $b$ between values $b_{0.5}^{H}$ and $b_{0.5}^{Hom}$
@@ 6. Singularities and Canards
As we have seen in case (i) of Section 5, when the system’s equilibrium is also a fold point, a subcritical Hopf bifurcation occurs, giving rise to a stable periodic solution (cf. Figure 6). In Sections 3 and 4, we found another periodic solution which is not related to the Hopf bifurcation. Thus, for different parameters, we have two stable periodic solutions with distinct dynamical properties. In this section, the canard phenomenon bridges the local periodic solution from the Hopf bifurcation to the global limit cycle.
As illustrated in Figure 6(d), the existence in (23) of a periodic solution around the equilibrium point implies that the orbit traverses a part of its trajectory close to the repelling region of the critical manifold $\mathcal{C}_{0}$. This behaviour seems to be contradictory since we found a solution that is close to the unstable branch of $\mathcal{C}_{0}$.
@@
Given a singular $(1,1)$–fast-slow system parametrised by $\lambda\in{\mathbb{R}}$. Let $\mathfrak{p}=\left(x,y\right)$ be a fold point. The point $\mathfrak{p}$ is called a singular fold if:
@@
Theorem 4 (Krupa & Szmolyan, 2001a, 2001b [22, 23]. See also Theorems 8.1.3 and 8.2.1 of [25]).
@@
$\lambda_{c}\left(\sqrt{\varepsilon}\right)=-\left(B+A\right)\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
@@
$\lambda_{\mathrm{H}}\left(\sqrt{\varepsilon}\right)=-B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
@@
An important consideration is the empirical difficulty in finding canards, since the smaller $\varepsilon$ is, the narrower the interval of $\lambda$ values for which canards appear. This small interval separates the $\lambda$ values at which a limit cycle occurs from those where it either undergoes a Hopf bifurcation or does not have any periodic solution. Figure 6(a) and its magnified Figures 6(b) and 6(c) illustrate this phenomenon. For this reason, the dynamics produced in the system around this small interval is called the canard explosion.
@@
$\bar{\mathrm{c}}_{\mathrm{H}}\left(\sqrt{\varepsilon}\right)=0\cdot\,\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right)$
@@
$\bar{\mathrm{c}}_{c}\left(\sqrt{\varepsilon}\right)=-\dfrac{3}{8}\,\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right).\qquad\diamond$
Example 6.2.
Now consider the case (ii) of Section 5 (with $c=0$) and the equilibrium $E_{+}=(x_{+},x_{+}/b)$, with $x_{+}=\sqrt{4-b^{-1}}$. Then $E_{+}$ is a regular singular fold point for $b=3/8$ with $x_{+}=4/3$ and there is a Hopf bifurcation point at $b=b_{\varepsilon}^{H}$ with $b_{\varepsilon}^{H}\to 3/8$ as $\varepsilon\to 0$. In order to apply Theorem 4 we change variables and parameters by
$(\bar{x},\bar{y},\lambda)=\left(x-x_{+},y-8x_{+}/3,b-3/8\right)$
to obtain the equivalent system
(29) $\begin{array}[]{lcl}\bar{x}^{\prime}&=&-\bar{y}-\bar{x}^{3}-3\bar{x}^{2}x_{+}\\ \bar{y}^{\prime}&=&\varepsilon\left(\bar{x}-\lambda\bar{y}-\frac{3}{8}\bar{y}\right).\end{array}$
In the notation of Theorem 4 we have
$l_{1}=-1\qquad l_{2}=-3x_{+}-\bar{x}\qquad l_{3}=0\qquad l_{4}=1\qquad l_{5}=1\qquad l_{6}=-\frac{3}{8}$
hence $A=-\dfrac{1}{8}\left(3+\dfrac{3}{4}\right)<0$ therefore the bifurcation is supercritical for $\lambda$, and also for $b$ in the original system (24). Since for $b<b_{\varepsilon}^{H}$ the eigenvalues of the matrix $J_{E_{+}}$ are positive, then the bifurcating periodic solution is unstable.
Figure 10 illustrates the emergence of canards in system (23) with $\varepsilon=0.5$ and $\varepsilon=0.1$. As we may see, the stable periodic solutions emerge from the subcritical Hopf bifurcation for $c=2/\sqrt{3}$ (analogous for $c=-2/\sqrt{3}$). As mentioned above, some of these periodic solutions remain close to the repelling region of the critical manifold, making them canards. In Figures 10(a) and 10(c), two types of orbits can be distinguished: the first type, coloured in yellow, consists of slow trajectories near $\mathcal{C}_{0\text{M}}$ and $\mathcal{C}_{0\text{R}}$ and one fast trajectory; the second type, coloured in green, which surrounds the other fold point, consists of three slow trajectories near $\mathcal{C}_{0\text{L}}$, $\mathcal{C}_{0\text{M}}$ and $\mathcal{C}_{0\text{R}}$ and two fast trajectories. We refer to these two types of canards as headless canard and headed canard, respectively [^2] This terminology is motivated by the literal meaning of the word “canard” in French (canard = duck). In this system, it is a bit more complicated to understand why, but if the reader tilts their head to the left they will see that the canard orbits resemble the shape of a duck, with or without head..
Note that the length of the canards grows almost instantaneously. As stated in [25], this was expected, but it is still surprising. The nearly vertical lines observed in Figures 10(b) and 10(d) show canards with lengths, $\mathcal{A}\left(c\right)\in\left[2,20\right]$ for $c\approx 1.150077$ in the case $\varepsilon=0.5$ and for $c\approx 1.153794$ in the case $\varepsilon=0.1$.
Figure 10. (a) - Periodic solutions of the system (23) for several values of $c$, with $\varepsilon=0.5$. Critical manifold in red, attracting regions (solid line) and repelling region (dashed line), fold points (black), canards without head (yellow), canards with head (green), and limit cycles (blue). (b) - Evolution of the length $\mathcal{A}$ of the canards of (a) as function of $c$. Unstable equilibrium point (dashed line), Hopf bifurcation point (black dot), length of canards without head and Hopf orbits (yellow), length of canards with head orbits (green), and length of limit cycles (blue). (c) - Periodic solutions of the system (23) for several values of $c$, with §$\varepsilon=0.1$ (conventions as in (a)) (d) - Evolution of the length of the canards of (c) as a function of the parameter $c$, conventions as in (b).
@@ 7. Discussion and Further Work
This article studies the dynamics of a fast-slow system derived from the FH-N model according to the geometric singular perturbation theory and the bifurcation theory methods. In particular, we provide an analytic proof that the fast-slow FH-N system presents relaxation oscillation dynamics as well as periodic solutions induced by Hopf bifurcation. We emphasise the emergence of a homoclinic orbit and canards connecting these two distinct dynamics. We illustrate each result with numerical computations. In each section, we have complemented the discussion of observed dynamics with stability and bifurcation analysis. This approach empowers the readers to navigate the parameter space, enabling them to find specific dynamical regimes of interest.
The FH-N equation is an excellent candidate to serve as a starting point for developing new methods to analyse transient and non-reciprocal coupled interactions. Cardiac and neuronal examples, in particular, highlight the FH-N model’s adaptability and significance in these areas [2].
When coupling two FH-N systems through the slow equations, we have performed numerical simulations that suggest the existence of mixed-mode oscillations where the dynamics is characterised by the existence of orbits that alternate periodically between large and small amplitude oscillations [13]. Recently, [21] showed that the coupling through the fast equations of two distinct FH-N models produces mixed-mode oscillations induced by a cusp singularity present in this system. We believe that a similar result may explain our numerical findings.
@@
$\phantom{\varepsilon\,}\dfrac{\mathrm{d}\theta}{\mathrm{d}t}=\omega,$
@@ Acknowledgments
The first and second authors were partially supported by CMUP, member of LASI, which is financed by national funds through FCT – Fundação para a Ciência e a Tecnologia, I.P., under the projects with reference UIDB/00144/2020 and UIDP/00144/2020. The third author has been supported by the Project CEMAPRE/REM – UIDB/05069/2020 financed by FCT/MCTES through national funds.